1.1 Prove that the characteristic polynomial of a 2 × 2 matrix A can be expressed as

λ^2 − tr(A)λ + det(A), where tr(A) is the trace of A.


1
Expert's answer
2022-04-23T09:51:10-0400

Let "A=\begin{pmatrix} a&b\\c&d \end{pmatrix}" . Then "\det (A)=ad-bc" and "\text{tr}(A)=a+d" .


The characteristic polynomial: "\det(A-\lambda I)=\begin{vmatrix}a-\lambda &b\\c&d-\lambda
\end{vmatrix}=(a-\lambda)(d-\lambda )-bc=\lambda^2-(a+d)\lambda +ad-bc=\lambda^2-\text{tr}(A)\lambda +\det(A)"